A 2-kg mass hangs at the end of a spring whose constant is k = 400 N/m. The mass is displaced (-) a
distance of 12cm and released. What is the acceleration at the instant the displacement is x = +7 cm?
a = (k / m) A or a = k x / m
1
Expert's answer
2020-12-10T11:02:09-0500

We can find the acceleration from the Hooke's Law:


"F=ma=-kx,""a=-\dfrac{kx}{m},""a=-\dfrac{400\ \dfrac{N}{m}\cdot 0.07\ m}{2\ kg}=-14\ \dfrac{m}{s^2}."


The sign minus means that when the displacement is +7 cm (downward) the acceleration is directed upward (hence, in the opposite direction to the motion of the 2-kg mass).

Answer:

"a=-14\ \dfrac{m}{s^2}," upward.